OpenStudy (anonymous):

A 1600kg car is traveling along a straight road at 20m/s . Two seconds later its speed is 24m/s .

OpenStudy (anonymous):

Ok. What's the rest of the question?

OpenStudy (anonymous):

the acceleration is 24-20/2 = 2m/s^2 The force applied is constant, and it is 1600 * 2 = 3200N The work the car used is F*d. d = V(initial) * t + 1/2 at^2 = 44, so work is 3200 * 44 = 140800 J, or 140 kJ

OpenStudy (anonymous):

That's probably one of the thing you're looking for right? :)